A number is 5 times another number. By adding 8 to each number ,the first number becomes only 3 times the second. What are the two numbers?
Answer by macston(5194)

.
F=first number=5S; S=second number
.
F+8=3(S+8)
F+8=3S+24 Substitute for F
5S+8=3S+24
5S=3S+16
2S=16
S=8 ANSWER 1: The second number is 8.
F=5S=5(9)=45 ANSWER 2: The first number is 40.
.
CHECK:
F+8=3(S+8)
40+8=3(8+8)
48=3(16)
48=48